OpenStudy (anonymous):

doreen uses 13 apples for every 2 pies she bakes how many apples will she use for twelve pies

OpenStudy (anonymous):

13 * 6

OpenStudy (anonymous):

13/2=6.5 (for every pie) so 6.5 x 12=78 apples

OpenStudy (anonymous):

i figured it out thanks

OpenStudy (anonymous):

For 2 pies 13 apples therefore for 12 pies 13x12/2 = 78